A typical cellular phone consumes an average of about 1.58 W of electrical power and operates on 3.80 V.

(a) What average current (in A) does the phone draw from its battery?

(b) Calculate the energy (in J) stored in a fully charged battery if the phone requires charging after 4.63 hours of use.
Answer #1

(a) as power is the product of voltage and current.

P = V * I

where P is average power , V is voltage and I is the current.

using the given values,

1.58 w = 3.80 v * I

solving for I,

I = 0.42 Ampere.

so the average current the battery draws is 0.42 Amp.

(b) as work is also a form of energy so,

E = P * t

where E is energy in joules,P is power in watt and t is the time in seconds.

using the known values,

E = 1.58 * (4.63 * 60 * 60)

E = 26335.4 joules

the energy stored in a fully charged battery is 26335.4 joules.